OpenStudy (anonymous):

Show that if \(a\) and \(b\) are positive integers, then \[\qquad \qquad \large\left(a + \frac12\right)^n +\left(b+ \frac12\right)^n\]is an integer for only finitely many positive integers \(n\).

OpenStudy (anonymous):

Still practicing, eh?

OpenStudy (anonymous):

Yeah. :/

OpenStudy (anonymous):

I have no idea.......(thinks binomial theorem, brain siezes up)

OpenStudy (anonymous):

How about this?\[\frac{(2a+1)^n + (2b+1)^n}{2^n}\]

OpenStudy (anonymous):

Hmm Binomial, nice thought.

OpenStudy (anonymous):

\[\frac{C_0 + C_12a + \ldots C_n(2a)^n + C_0 + C_12b + \ldots +C_n(2b)^n}{2^n}\]

OpenStudy (anonymous):

Or, maybe this\[(1+2a)^n + (1+2b)^n = 2^n k, \quad k \in \mathbb{Z}\]

OpenStudy (anonymous):

\[2C_0 + 2C_1(a+b) + 4C_2(a^2+b^2) + \ldots +2^nC_n(a^n+b^n)=2^n k \]

OpenStudy (anonymous):

I don't think it's gonna help :/

OpenStudy (anonymous):

Do you know how to factorize \(x^n + y^n\)?

OpenStudy (anonymous):

Hang on a minute, if the result has to be integer how does the 1/2 at the end play out? (a+b)^n always has b^n at the end

OpenStudy (anonymous):

Some other terms have to make it up to an integer.

OpenStudy (anonymous):

Not the first term and all the other terms are combinations

OpenStudy (anonymous):

And the numerator must be even (per your equation above)

OpenStudy (anonymous):

I think I know 2k +1 is always an Odd term. \[\frac{(2a+1)^n + (2b+1)^n}{2^n}\] Odd^n + Odd^n = Even and Even%2^n = 0, what kind of even numbers have remainder Zero upon division by 2?

OpenStudy (anonymous):

For 2^1, every value for a and b should work

OpenStudy (anonymous):

For 2^2 or 4, there must be less values.

OpenStudy (anonymous):

My head is hurting....:-(

OpenStudy (anonymous):

I will try to come back to it later......

OpenStudy (anonymous):

Sure, and thanks for trying :-)

OpenStudy (anonymous):

One of us under the alias of ManiSarkarCallisto posted this question in M.SE : http://math.stackexchange.com/questions/139035/

OpenStudy (anonymous):

Lol, that was me Lololololol

OpenStudy (anonymous):

:-)